# Improving Visual Place Recognition with Sequence-Matching Receptiveness Prediction

### v1

@@ Abstract
In visual place recognition (VPR), filtering and sequence-based matching approaches can improve performance by integrating temporal information across image sequences, especially in challenging conditions. While these methods are commonly applied, their effects on system behavior can be unpredictable and can actually make performance worse in certain situations. In this work, we present a new supervised learning approach that learns to predict the per-frame sequence matching receptiveness (SMR) of VPR techniques, enabling the system to selectively decide when to trust the output of a sequence matching system. The approach is agnostic to the underlying VPR technique. Our approach predicts SMR—and hence significantly improves VPR performance—across a large range of state-of-the-art and classical VPR techniques (namely CosPlace, MixVPR, EigenPlaces, SALAD, AP-GeM, NetVLAD and SAD), and across three benchmark VPR datasets (Nordland, Oxford RobotCar, and SFU-Mountain). We also provide insights into a complementary approach that uses the predictor to replace discarded matches, as well as ablation studies, including an analysis of the interactions between our SMR predictor and the selected sequence length. We will release our code upon acceptance.
@@ I Introduction
One of the key challenges faced by all classes of VPR Systems is environmental appearance change, caused by any of a number of factors – lighting conditions, seasons, day-night cycles, and structural change [9, 10]. One of the common methods for improving the performance of VPR systems to such challenges is to filter input over time and space, for example through sequence matching techniques [11, 12, 13, 14, 15, 16, 17]. But sequence matching is not a perfect solution: sometimes using sequence matching can degrade performance, depending on factors like varying platform velocities or path divergence [18]. Knowing when to use and trust sequence matching as a VPR performance enhancement technique is the focus of the research presented in this paper.
We present techniques for training a system to predict whether the result of applying sequence matching is likely to yield an improved place recognition output or not, and utilize this by using this predictor to selectively decide whether to trust the output of a sequence matcher (Figure 1). Because our method is focused on assessing the performance contribution of the sequence matching specifically, it is agnostic to the underlying VPR technique. This general applicability is demonstrated in the extensive experimental evaluation, whereby we evaluate our system across seven VPR techniques, including both recent state-of-the-art methods and classical approaches, and three benchmark datasets. These experiments demonstrate that our predictor, even though not perfect, is able to substantially improve average VPR performance across this extensive range of dataset and method combinations.
@@
Whilst introspection for VPR is a fledgling but growing field [19, 20, 21], all prior work has primarily focused on assessment at the instantaneous match level: our contribution here is the first to apply to the widely relevant scenario where filtering or sequence matching is applied.
The paper proceeds as follows. Section II reviews related work across the fields of visual place recognition, sequence matching, and introspection. Section III details our methodology, including how we process the underlying VPR system’s output to train and apply a match quality predictor, and how that prediction process is used to remove (and potentially restore) VPR matches. The experimental setup across seven methods and three benchmark datasets is covered in Section IV, with the overall performance and ablation and analysis results provided in Section V. Finally, the paper concludes in Section VI with discussion and identification of promising areas of future work.
@@ II-A Visual Place Recognition
Many VPR approaches use an image retrieval approach which involves extracting meaningful feature representations from both reference and query images [9]. The goal is for the true reference image to be the best match for a given query image, despite challenges caused by appearance, viewpoint, and perceptual aliasing [18].
Early VPR methods relied on hand-crafted feature representations [22], which later evolved into deep learning representations to improve feature robustness [23]. More recent VPR methods demonstrate even greater robustness to challenging conditions [24, 25, 26, 27, 28, 29], which is enabled by the use of Transformer-based architectures, pre-trained or fine-tuned foundation models [29, 27, 28], feature fusion techniques [26], and training over multiple image groups [24, 25]. Despite these great advances in VPR literature, these VPR methods can still be susceptible to sudden environmental changes, the appearance of dynamic objects, and perceptual aliasing, which can lead to incorrect matches.
@@ II-B Sequence Matching for Visual Place Recognition
Sequence matching and filtering are common strategies that are widely applied to improve the robustness of single-frame VPR methods. These techniques use the existing temporal information in scenarios where VPR is typically applied, e.g., when a mobile robot moves along a path [18].
Sequence matching approaches for VPR include the following. Similarity-based techniques are applied as a post-processing step to the similarity or distance scores obtained from the matching process [18, 11, 12, 13] such as SeqSLAM [11]. These approaches aggregate matching scores from reference and query subsets, defined by a predefined sequence length [18]. Feature-based techniques aggregate consecutive reference and query image features into a single feature representing the entire sequence [16, 17]. Learning-based approaches train sequential feature descriptors to capture temporal information, such as in SeqMatchNet [14] and deep-learning or transformer-based methods [15].
These sequence matching approaches take advantage of temporal data, which can potentially be more compute-intensive and have higher latency compared to single-frame VPR techniques. Most importantly to the context of this paper, while these methods often improve overall performance, there is typically no self-characterization whether the sequence matcher improves performance on a match-by-match basis.
@@ II-C Introspection for Visual Place Recognition
Localization integrity has been explored in GNSS navigation systems [31], and introspection of autonomous systems [32, 33]. Recently, in the context of VPR, researchers have explored probabilistic approaches to estimate uncertainty and predict the integrity of single-frame VPR systems, both on benchmark datasets [19] and in real-time robotic applications [20] to improve their matching outcomes. Another VPR work [21] predicted matches within ground-truth tolerance and weighted single-frame match scores, considered as most-likely true matches, to make the VPR system more responsive to sequence matching.
Similar to the works above [19, 20, 21], in this work, we consider predicting the likelihood that the best match is a true match. However, while these works focus on predicting or improving single-frame VPR match outcomes, this work focuses on predicting whether a single-frame VPR technique, on a per-frame basis, would benefit from sequence matching. Unlike most SLAM methods, which do most of their assessment of how to integrate sensory and motion information over time by varying the weighting of these input modalities, our method assesses the likely quality of the proposed matching output after this integration has already occurred.
@@ III Methodology
Our approach focuses on predicting sequence matching receptiveness (SMR) on a per-frame basis, which is agnostic to the underlying VPR model. The goal is to predict the impact of applying sequence matching to a query prediction: would the prediction remain unchanged; would an incorrect prediction be corrected; or finally, would correct prediction turn into an incorrect one?
We frame this as a classification task with a one-to-one correspondence between reference and query images. Our approach for predicting sequence matching receptiveness includes extracting attributes111We refer to features for our SMR predictor as attributes to distinguish them from image features extracted by VPR models. from the distance scores of a query image, and its previous frames, and feeding them into a supervised Multi-layer Perceptron (MLP) classifier to predict the query’s outcome after sequence matching.
@@ III-A Distance Matrix
This section details the construction of distance matrices from single-frame VPR techniques, which are used for extracting classifier attributes in Section III-C. Given a set of reference images and query images, we construct an distance matrix , where each element represents the distance scores between the image feature representations of the -th reference image and the -th query image. Common algorithms used for calculating these distances include Euclidean and Cosine distances. Our SMR prediction model relies on the distance scores of a query frame and its previous frames to all reference images obtained from the single-frame VPR technique. This is essentially a vertical slice of a distance matrix, and its width is determined by the selected sequence matching length .
@@ III-C SMR Predictor Attributes
This section describes the per-frame attributes extracted from the distance scores of a VPR method. We define as the distance scores matrix, which is a slice of containing scores for the current query frame and its previous frames. We compute attributes based on the diagonal entries matrix, . These entries, spanning from the top-left to bottom-right, capture sequential correspondences between the query and reference images over the sequence length. Specifically, for each row of diagonal entries , is defined as:
@@ III-C1 Minimum Sum Rate
To capture sequential alignment consistency between the query’s previous frames, we compute the ratio between the sum of horizontal entries (with the top lowest row-wise mean) to the sum of diagonal entries (with the top lowest diagonal-wise mean) from the distance scores matrix, . For horizontal entries in row , the mean distance score across the previous frames is computed as:
@@
where is the row of diagonal entries with the top lowest mean from , is the row of horizontal entries with the top lowest mean from mean distance scores vector, and is a small constant to prevent division by zero.
@@ III-C2 Minimum Value Rate
To capture the reliability of the strongest match, we compute the ratio between the top lowest distance scores for the query and the distance score for the query at the row of diagonal entries with the lowest top sum.
where rank selects the top row with the lowest distance score for the query, and is the row of diagonal entries with the top lowest distance scores sum.
III-C3 Global Block Sum Ratio
@@
Finally, we compute the global block sum ratio as:
III-C4 Global Group Sum Ratio
Additionally, we compute the ratio between the sum of the diagonal entries for the row with the top lowest diagonal sum, and the corresponding sum of the diagonal group entries in the same row. We define a diagonal group by averaging the values of the diagonal entries in a window centered around each row. Specifically, for each row, we consider the rows before and after it, and compute the mean of the values within this window. This process smooths the diagonal by incorporating surrounding context. We define a group as:
Finally, we compute the global group sum ratio as:
@@ III-C5 Final Feature Representation
We compute these attributes for each frame based on the top lowest entries, which are ranked by their sums. Hence, we obtain a attributes vector for each frame. Noting that, we train and evaluate our MLP-based SMR predictor based only on the top attributes for each frame.
@@ III-D SMR Predictor Supervised Learning Method
: Correct before and incorrect after sequence matching.
: Correct before and after sequence matching.
For a query frame , the top attributes are , with output .
@@ III-E Match Removal
Leveraging the predictor to improve performance involves removing initial proposed matches that are predicted as “incorrect after sequence matching” by our predictor (). We note here that whilst ideally the prediction would be perfect, in the results section we show that even an imperfect predictor is able to produce a significant overall improvement in VPR performance. By using our predictor to remove some matches, the aim is to primarily remove actual false positives, whilst minimizing the number of true positives that are removed (which makes them into false negatives).
@@ III-F Match Restoration
If matches can be removed because they are unlikely wrong, can we promote other highly ranked matches to replace them? That is the aim of this investigation, where we use our predictor on the next top ranked matches (as ranked by the distance scores of the VPR system), and replace the “removed” false positive with the top ranked match that has the highest score from the predictor system. The aim of such an approach is to provide a mechanism for increasing recall, whilst minimizing the drop in precision – the inverse of our main contribution. We provide this analysis here not as part of our main methodological contribution, which already by itself improves VPR performance across many methods and datasets without this latter step – but rather to provide some initial insights into this natural follow-on question.
@@ IV-A Implementation Details
We normalize the distance scores matrix, , by dividing each entry by the maximum absolute value. As the number of query samples for each of the output label classes varies across all samples, we used the Synthetic Minority Over-sampling Technique (SMOTE) [35] to upsample the minority classes for training only. We used for the Global Group Sum Ratio attribute. For the MLP classifier, we use three hidden layers of 128 neurons each, with the Rectified Linear Units (ReLU) activation and Adam optimization. The initial learning rate is set to , and . We employ Stratified K-Fold cross-validation and use the macro F1-score for evaluation. We implemented our MLP classifier using the SkLearn Python library [36] on a machine with an i7 CPU, RTX 3080 GPU.
@@ IV-C Datasets
The Nordland dataset [38]: This 728 km train route in Norway spans across four seasons: spring, summer, fall, and winter. Segments with speeds below 15 km/h were excluded as done in [39]. We used the summer traverse as the reference and the winter traverse as the query data. We used reference and corresponding query images for training, geographically separate images for validation, and for testing.
The Oxford RobotCar dataset [40]: This dataset includes over 100 traverses in Oxford under various conditions, such as different times of day and seasons. We used the front left stereo frames from the Rain traverse as the reference and from the Dusk traverse as the query222Rain: 2015-10-29-12-18-17, Dusk: 2014-11-21-16-07-03 as in [41]. We used reference and query images for training, geographically separate images for validation, and for testing.
SFU Mountain dataset [42]: This dataset with 385 frames was recorded by a mobile robot in a semi-structured bushland terrain at Burnaby Mountain, British Columbia, Canada, under challenging appearance conditions. As in [43], we used the right stereo frames and used the dry traverse as the reference and the dusk traverse as the query. We used the first reference and query images for training, geographically separate images for validation, and for testing.
@@ IV-D Performance Metrics
| VPR+SM | 0.2525 | 0.0575 | 0.1584 | 0.0287 | 0.0808 | 0.2210 | 0.0673 | 0.1237 | ||
| Nordland | VPR+SM+Pred | 0.1752 | 0.0552 | 0.0751 | 0.0284 | 0.0559 | 0.0791 | 0.0037 | 0.0675 | |
| Reduction (%) | 30.62 | 3.96 | 52.58 | 0.95 | 30.79 | 64.21 | 94.50 | 39.66 | 7 / 7 | |
@@ V-A Performance of our SMR Predictor
In 14 out of 21 evaluated VPR method and dataset configurations, using our predictor results in an AOC reduction compared to the VPR system with sequence matching. In 4 out of 21 cases, all from the SFU Mountain dataset, the AUC-PR of the VPR system with sequence matching is already capped at 100%. We show that applying our predictor maintains this perfect performance, ensuring no degradation in the VPR+SM system’s performance. In 2 out of 21 cases, both from the Oxford RobotCar dataset, there is a slight performance degradation, with the AOC increasing from to for AP-GeM and from to for CosPlace.

### v2

@@ Abstract
In visual place recognition (VPR), filtering and sequence-based matching approaches can improve performance by integrating temporal information across image sequences, especially in challenging conditions. While these methods are commonly applied, their effects on system behavior can be unpredictable and can actually make performance worse in certain situations. In this work, we present a new supervised learning approach that learns to predict the per-frame sequence matching receptiveness (SMR) of VPR techniques, enabling the system to selectively decide when to trust the output of a sequence matching system. Our approach is agnostic to the underlying VPR technique and effectively predicts SMR, and hence significantly improves VPR performance across a large range of state-of-the-art and classical VPR techniques (namely CosPlace, MixVPR, EigenPlaces, SALAD, AP-GeM, NetVLAD and SAD), and across three benchmark VPR datasets (Nordland, Oxford RobotCar, and SFU-Mountain). We also provide insights into a complementary approach that uses the predictor to replace discarded matches, and present ablation studies including an analysis of the interactions between our SMR predictor and the selected sequence length.
@@ I Introduction
One of the key challenges faced by all classes of VPR systems is environmental appearance change, caused by any of a number of factors: lighting conditions, seasons, day-night cycles, and structural change [9, 10]. One of the common methods for improving the performance of VPR systems in challenging conditions is to filter input over time and space, for example, through sequence matching techniques [11, 12, 13, 14, 15, 16, 17]. But sequence matching is not a perfect solution: sometimes using sequence matching can degrade performance, depending on factors like varying platform velocities or path divergence [18]. Knowing when to use and trust sequence matching as a VPR performance enhancement technique is the focus of the research presented in this paper.
We present techniques for training a system to predict whether the result of applying sequence matching improves the likelihood of the match being correct, and use this predictor to selectively decide whether to trust the output of a sequence matcher (Figure 1). As our method is focused on assessing the performance contribution of the sequence matching specifically, it is agnostic to the underlying VPR technique. This general applicability is demonstrated in extensive experimental evaluations, where we evaluate our system across seven VPR techniques, including both recent state-of-the-art methods and classical approaches, and three benchmark datasets. These experiments demonstrate that our predictor, even though not perfect, is able to substantially improve average VPR performance across this range of dataset and method combinations.
@@
Whilst introspection for VPR is a fledgling but growing field, all prior work [19, 20, 21] has primarily focused on assessment at the instantaneous match level: our contribution here is the first to apply to the widely relevant scenario where filtering or sequence matching is applied.
The paper proceeds as follows. Section II reviews related work in visual place recognition, sequence matching, and introspection. Section III details our methodology, including processing the VPR system’s output to train and apply a match quality predictor, and using predictions to remove (and potentially restore) VPR matches. Section IV covers the experimental setup across seven methods and three benchmark datasets. Section V presents overall performance, ablations, and analysis. Section VI concludes with discussion and future work. Our code is available online111https://github.com/QVPR/vpr-smr-predictor.
@@ II-A Visual Place Recognition
Many VPR approaches use an image retrieval approach which involves extracting meaningful feature representations from both reference and query images [9]. The goal is for the true reference image to be the match with the highest similarity for a given query image, despite challenges caused by appearance, viewpoint, and perceptual aliasing [18].
Early VPR methods relied on hand-crafted feature representations [22], which later evolved into deep learning representations to improve feature robustness [23]. More recent VPR methods demonstrate even greater robustness to challenging conditions [24, 25, 26, 27, 28, 29], which is enabled by the use of Transformer-based architectures, pre-trained or fine-tuned foundation models [29, 27, 28], feature fusion techniques [26], and training over multiple image groups [24, 25]. However, these VPR methods can still be susceptible to sudden environmental changes, the appearance of dynamic objects, and perceptual aliasing, which can lead to incorrect matches.
@@ II-B Sequence Matching for Visual Place Recognition
Sequence matching and filtering are common strategies that are widely applied to improve the robustness of single-frame VPR methods. These techniques use the existing temporal information in scenarios where VPR is typically applied, e.g. when a mobile robot moves along a path [18].
Similarity-based sequence matching techniques are applied as a post-processing step to the similarity or distance scores obtained from the matching process [18, 11, 12, 13] such as SeqSLAM [11]. These approaches aggregate matching scores from reference and query subsets, defined by a predefined sequence length [18]. Feature-based techniques aggregate consecutive reference and query image features into a single feature representing the entire sequence [16, 17]. Learning-based approaches train sequential feature descriptors to capture temporal information, such as in SeqMatchNet [14] and deep-learning or transformer-based methods [15].
These sequence matching approaches take advantage of temporal data, which can potentially be more compute-intensive and have higher latency compared to single-frame VPR techniques. Most importantly to the context of this paper, while these methods often improve overall performance, there is typically no self-characterization of whether the sequence matcher improves performance on a match-by-match basis.
@@ II-C Introspection for Visual Place Recognition
Localization integrity has been explored in GNSS navigation systems [31], and introspection of autonomous systems [32, 33]. Recently, in the context of VPR, researchers have explored probabilistic approaches to estimate uncertainty and predict the integrity of single-frame systems, both on benchmark datasets [19] and in real-time robotics [20], to improve matching outcomes. Another study [21] predicted matches within ground-truth tolerance and weighted the scores of likely true single-frame matches to make the VPR system more responsive to sequence matching.
Similar to [19, 20, 21], we consider predicting the likelihood that the best match is a true match. While those works focus on predicting or improving single-frame VPR outcomes, our work predicts, on a per-frame basis, whether a single-frame VPR technique would benefit from sequence matching. Unlike most SLAM methods, which proactively fuse sensory and motion information over time, our method retrospectively assesses the likely quality of the proposed match after sequence integration has already occurred.
@@ III Methodology
Our approach focuses on predicting sequence matching receptiveness (SMR) on a per-frame basis, and is agnostic to the underlying VPR model. The goal is to predict the impact of applying sequence matching to a query prediction: would the prediction remain unchanged; would an incorrect prediction be corrected; or, finally, would a correct prediction turn into an incorrect one?
We frame this as a classification task with a one-to-one correspondence between reference and query images. Our approach predicts sequence matching receptiveness by extracting attributes from the distance scores of a query and its previous frames (with their number defined by the sequence length) and feeding them into a supervised Multi-layer Perceptron (MLP) classifier to predict the query’s outcome after sequence matching. We refer to features for our SMR predictor as attributes to distinguish them from image features extracted by VPR models.
@@ III-A Distance Matrix
Here, we detail how the distance matrices from single-frame VPR techniques are constructed, which are used for extracting classifier attributes in Section III-C. Given a set of reference and query images, we construct an distance matrix , where each element represents the distance scores between the image feature representations of the -th reference image and the -th query image. Common algorithms used for calculating these distances include Euclidean and Cosine distances. Our SMR prediction model relies on the distance scores of a query frame and its previous frames to all reference images obtained from the single-frame VPR technique. This is essentially a vertical slice of a distance matrix, and the number of previous frames is determined by the selected sequence matching length .
@@ III-C SMR Predictor Attributes
This section describes the per-frame attributes extracted from the distance scores of a VPR method. We define as the distance scores matrix, which is a slice of containing scores for the current query frame and its previous frames. We calculate the attributes based on the diagonal entries matrix, . These entries, spanning from the top-left to bottom-right of the distance scores matrix , capture sequential correspondences between the query and reference images over the sequence length. For each row of diagonal entries , of query is defined as:
@@ III-C1 Minimum Sum Rate
To capture sequential alignment consistency between the query’s previous frames, we compute the ratio between the sum of horizontal entries (with the top lowest row-wise mean) to the sum of diagonal entries (with the top lowest diagonal-wise mean) from the distance scores matrix, . For horizontal entries in row of query , the mean distance score across the previous frames is computed as:
@@
where is the row of diagonal entries with the top lowest mean from , is the row of horizontal entries with the top lowest mean from mean distance scores vector, and is added to prevent division by zero; alternatively, a max operation could be used.
@@ III-C2 Minimum Value Rate
To capture the reliability of the strongest match, we compute the ratio between the top lowest distance scores for the query and the distance score for the query at the row of diagonal entries with the lowest top sum:
where is the row of diagonal entries with the top lowest distance scores sum.
III-C3 Global Block Sum Rate
@@
Finally, we compute the global block sum rate as:
III-C4 Global Group Sum Rate
Finally, we compute the ratio between the sum of the diagonal entries for the row with the top lowest diagonal sum, and the corresponding sum of the diagonal group entries in the same row. We define a diagonal group by averaging the values of the diagonal entries in a window centered around each row. Specifically, for each row, we consider the rows before and after it, and compute the mean of the values within this window. This process smoothens the diagonal by incorporating surrounding context. We define a group as:
Finally, we compute the global group sum rate as:
@@ III-C5 Final Feature Representation
We compute these attributes for each frame based on the top lowest entries, which are ranked by their sums. Hence, for each query frame, we extract four attributes for each top lowest entries. Note, we train and evaluate our MLP-based SMR predictor based only on the top attributes.
@@ III-D SMR Predictor Supervised Learning Method
: Correct before and incorrect after sequence matching.
@@
: Incorrect before and after sequence matching.
@@ III-E Match Removal
Leveraging the predictor to improve performance involves removing initial proposed matches that are predicted as “incorrect after sequence matching” by our predictor (). We note here that whilst ideally the prediction would be perfect, in the results section we show that even an imperfect predictor is able to produce a significant overall improvement in VPR performance. By using our predictor to remove some matches, the aim is to primarily remove actual false positives, whilst minimizing the number of true positives that are removed (which turns them into false negatives).
@@ III-F Match Restoration
If matches can be removed because they are likely wrong, can we promote other highly ranked matches to replace them? That is the aim of this investigation, where we use our predictor on the next top ranked matches (as ranked by the distance scores of the VPR system), and replace the “removed” false positive with the top ranked match that has the highest score from the predictor system. The aim of such an approach is to provide a mechanism for increasing recall, whilst minimizing the drop in precision – the inverse of our main contribution. We provide this analysis here not as part of our main methodological contribution, which already by itself improves VPR performance across many methods and datasets without this latter step – but rather to provide some initial insights into this natural follow-on question.
@@ IV-A Implementation Details
We normalize the distance score matrix, , by dividing each entry by the maximum absolute value in . As the number of query samples per output class varies, we used the Synthetic Minority Over-sampling Technique (SMOTE) [35] to upsample minority classes for training only. We use for the Global Group Sum Ratio attribute. Our MLP classifier has three hidden layers of 128 neurons each, with the Rectified Linear Units (ReLU) activation and Adam optimization. The initial learning rate is set to , and . We apply Stratified K-Fold cross-validation and evaluate using the macro F1-score. We implemented our classifier using the SkLearn Python library [36] on a machine with an i7 CPU, RTX 3080 GPU.
@@ IV-C Datasets
The Nordland dataset [38]: This 728 km train route in Norway spans across four seasons: spring, summer, fall, and winter. Segments with speeds below 15 km/h were excluded as done in [39]. We used the summer traverse as reference and winter as query. We used reference and corresponding query images for training, geographically separate images for validation, and for testing.
The Oxford RobotCar dataset [40]: This dataset includes over 100 traverses in Oxford under various conditions, such as different times of day and seasons. We used the front left stereo frames from the Rain traverse, 2015-10-29-12-18-17, as reference and from the Dusk traverse, 2014-11-21-16-07-03, as query as in [41]. We used reference and query images for training, geographically separate images for validation, and for testing.
SFU Mountain dataset [42]: This dataset with 385 frames was recorded by a mobile robot in a semi-structured bushland terrain at Burnaby Mountain, British Columbia, Canada, under challenging appearance conditions. As in [43], we used the right stereo frames and used the dry traverse as reference and the dusk traverse as query. We used the first reference and query images for training, geographically separate images for validation, and for testing.
@@ IV-D Performance Metrics
| VPR+SM | 0.6522 | 0.0220 | 0.0155 | 0.0061 | 0.2478 | 0.2164 | 0.0017 | 0.1660 | ||
| Nordland | VPR+SM+Pred | 0.4862 | 0.0136 | 0.0106 | 0.0042 | 0.1274 | 0.1050 | 0.0011 | 0.1069 | |
| Reduction (%) | 25.46 | 38.17 | 31.31 | 31.96 | 48.61 | 51.46 | 37.55 | 37.79 | 7 / 7 | |
@@ V-A Performance of our SMR Predictor
In 15 out of 21 evaluated VPR method and dataset configurations, using our predictor results in an AOC reduction compared to the VPR system with sequence matching. In 4 out of 21 cases, all from the SFU Mountain dataset, the AUC-PR of the VPR system with sequence matching is already capped at 100%. We show that applying our predictor maintains this perfect performance, ensuring no degradation in the VPR+SM system’s performance. In 2 out of 21 cases, both from the Oxford RobotCar dataset, there is a slight performance degradation, with the AOC increasing from to for AP-GeM and from to for CosPlace.